# references.bib

### v2

@@ 3.1. Common ground process
3.1.1   Step 1: Collective dialogue
The first step towards a collective dialogue is developing the planned prompts to which participants will respond and deciding on what demographics to gather from participants. Prompt drafting is informed by the goals of the dialogue: specific issues to find common ground on, areas to create more mutual understanding among participants, etc. Draft prompts are then reviewed and iterated until deemed acceptable by a stakeholder group with representation from all sides. The choice of demographics are determined by (a) demographic dimensions likely to manifest relevant divisions, (b) additional demographics needed to assess representativeness, and (c) constraints imposed by the need for statistical anonymity (bravohermsdorff2022statistical).
@@ 3.1.1   Step 1: Collective dialogue
Each collective dialogue generates the following raw data: a set of natural language statements $S$ responding to each question; a sparse sampling of participant votes on each statement $s\in S$; the demographics collected for each participant, and data for any other non-collective-response questions asked during the dialogue. Vote inference (konya2022elicitation) is then used to complete the agreement-vote matrix with elements $v(s,p)\in\{1,0\}\equiv\{agree,disagree\}$ that indicate the (stated or inferred) agreement of participant $p$ with statement $s$.
3.1.2   Step 2: Identifying bridging statements
@@
The goal of this step is to identify a set of “bridging” statements $B\subseteq S$ that are likely to represent points of common ground. To do this, we computed two bridging metrics for each statement $s\in S$: (1) max-min agreement $a_{\textsf{min}}(s,\mathcal{G})$ across defined demographic groups $\mathcal{G}$, and (2) the learned parameters $(f_{s},i_{s})$ from the Community Notes algorithm (see Section 2.2.2). As a first set of bridging statements, we took those statements $s\in S$ for which $a_{\textsf{min}}(s,\mathcal{G})$ was greater than a given threshold. This gives the set of statements that have the highest lowest agreement across the specified set $\mathcal{G}$ of demographic groups. However, when $G$ contains many demographic segmentations — ie. by age, gender, religiosity, etc — this approach will exclude a statement that may otherwise bridge a critical axis of division just because it is predicted (perhaps incorrectly) to have low agreement among one of the demographic sub-groups. To mitigate this risk, we created a second set of bridging statements using the Community Notes approach, consisting of those for which the Euclidean distance $d((f_{s},i_{s}),(0,2))$ was below a given threshold. This approach relies primarily on the note intercept $i_{s}$ as a bridging metric, but also penalizes statements for having non-zero slant, as indicated by the note factor $f_{s}$. The complete set of bridging statements $B$ is taken to be the union of the sets of statements emerging from both approaches.
3.1.3   Step 3: Distillation into collective statements
@@
The set of bridging statements identified in the previous step typically includes multiple restatements of the same idea, compound statements covering multiple ideas, and sub-optimal articulations of key ideas. The goal of Step 3 is to create well-articulated versions of each of the unique ideas that appear in the bridging statements in a way that preserves participants’ choices of specific words and phrases. We use an LLM to do this in two steps. First, the LLM is prompted to extract the unique ideas from the set of bridging statements. Next, the LLM is then prompted to generate well-articulated versions of each of those ideas in a way that preserves the specific words and phrases found in the original statements. Included in this prompt are the set of unique ideas, the text of the bridging statements from which the unique ideas were derived, and a set of exemplar statements that demonstrate the style in which collective statements should be written. Finally, the collective statements proposed by the LLM are then reviewed by local language experts for linguistic consistency with the original bridging statements, and iterated as needed until a final set of collective statements $C$ is deemed fair by a stakeholder group with representation from all sides. See Appendix A for LLM prompts used here.
3.1.4   Step 4: Final vote
@@
3.1.5   Joint reviews
Even though the final vote serves as a hedge against biases that could cause collective statements to not represent common ground, the existence or perception of unfair bias at any point in the process can still undermine legitimacy. To mitigate this risk, at key junctures in the process where bias might be introduced, joint reviews were conducted by representatives of each side before moving on to the next step (see Figure 1). Specifically, joint reviews were conducted on: the content and wording of all messages, questions, and collective response prompts that would be seen by participants in the collective dialogue, the collective statements that would go on to the final vote, and the results computed from the final vote data.
@@ 3.2.2   Strategic sequencing
What constitutes a “side”? It may seem obvious that “Israelis” and “Palestinians” would be the two sides. However, this binary approach would have risked marginalizing Palestinian citizens of Israel, may not fully identify with either Israelis or Palestinians. To accommodate this, we chose to treat Palestinian citizens of Israel as a third independent group, along side Israeli Jews and Palestinians from the West Bank and Gaza. In the remainder of this paper, we use $\mathcal{G}$ to denote this set of groups, namely
@@
During the uninational phase, the group that went first needed to do so without the benefit of any trust-building feedback from earlier cycles. Given the existing power asymmetry, it was decided that Israeli Jews would thus go first. As the second major group, Palestinians from the West Bank and Gaza went second, followed by the Palestinian citizens of Israel. Finally, after these three uninational dialogues, we conducted a joint dialogue with participants from all groups to attempt to identify cross-border common ground (Figure 1).
@@ 3.3. Languages
It was critical for legitimacy and inclusivity that participants could fully engage in their primary languages. The uninational cycles would be mono-lingual (Hebrew or Palestinian Arabic), and the joint cycle would need to be multilingual. The main challenge for the mono-lingual cycles was the LLM pipeline used to generate collective statements. Early versions of the pipeline used in prior work (konya2023dem; konya2024chain) had proven effective when the prompts, examples, bridging statements, and generated outputs were all in English, but it was unclear how to approach Hebrew and Palestinian Arabic. Should prompts or examples still be in English? And who decides? To pragmatically address these questions we set up two micro-experiments. We compiled a set of plausible bridging statements in each language and asked native speakers with prompt engineering experience to translate the English prompts and examples. Different variations of the pipeline were created by swapping out different permutations of the English prompts and examples with translated versions. Collective statements were generated using each variation. Then, blind to how they were generated, native speakers were asked which set of collective statements best reflected the content, culture, and language of the original statements they were based on. In all cases, reviewers preferred the statements generated by a pipeline with prompts and examples in their native language, so these variants were used for the uninational phase.
@@ 3.4. Equal power metrics
For the joint cycle, it was critical that result metrics computed from the final vote data (including both the ranking and agreement votes) give each side equal influence even with unequal shares of participants. To that end, we computed three equal-power metrics — that is, metrics that are invariant to the portion of participants from each side — for each collective statement: max-min agreement, Dowdall score, and IRV rank.
  • Max-min agreement: Max-min agreement, the same metric used for bridging-based ranking in Step 2, was computed again using the final agreement votes. Specifically, for each collective statement $s$, max-min agreement was computed as $\min\{\alpha(s,G)\mid G\in\mathcal{G}\}$ where $\alpha(s,G)$ is the fraction of participants in group $G$ that voted “agree” or “strongly agree” on statement $s$. This metric inherently provides equal power across groups and offers a straightforward interpretation: it indicates the minimum level of agreement across all defined groups for statement $s$.
  • Dowdall score: Similar to a Borda count (brandt2016handbook), a Dowdall score is a cardinal metric computed from rank votes using a positional scoring rule that quantifies the strength of relative preference among a set of alternatives for a voting population (fraenkel2014borda). Let $\textsf{Position}(s,k;P)$ give the fraction of participants (in population $P$) who ranked statement $s$ in position $k$. A Dowdall score is computed as
@@
    where $\textsf{Points}(k)=1/k$ corresponds to the ‘points’ assigned to each position. We adapt this to be an equal power metric by taking the average of Dowdall scores computed independently for each side:
@@
  • IRV rank: An instant runoff voting (IRV) rank is an ordinal metric derived from rank votes by iteratively eliminating the option with the fewest first-place votes among remaining options (brandt2016handbook). Let $S_{n}$ be the set of statements remaining in round $n$, and $\textsf{First}(s,n;P)$ give the fraction of participants (in some population $P$) who rank statement $s$ first among $S_{n}$. The statements remaining in round $n+1$ are $S_{n+1}=S_{n}\setminus\{\arg\min_{s\in S_{n}}\textsf{First}(s,n;P)\}$. [^5] This formulation assumes no ties for last place, which was true in practice with our data. Denoting $n_{s}$ to be the round where statement $s$ was eliminated, and $|C|$ the total number of statements being ranked, the IRV rank is $\textsf{IRV}(s)=|C|-n_{s}+1$. To make this an equal power metric, the ‘first place fraction’ $\textsf{First}(s,n)$ used to eliminate statements in each round is computed as an average of the first place fractions within each group. That is,
@@ 4. Results
The interpretation of the results presented herein necessitates careful consideration of several contextual factors. First, given the sensitive nature of the issues addressed in these dialogues and their direct relevance to ongoing peacebuilding efforts, the results shared are intentionally selective. Specifically, we present: (1) findings from the uninational phase that demonstrably facilitated the subsequent joint phase, and (2) collective statements of common ground from the joint phase that were explicitly intended for external communication. Furthermore, the quantitative results we share are limited to metrics that give equal power representation across primary participant groups.
Second, we emphasize that these findings should not be understood as the views of the broader Israeli and Palestinian populations. The participant pool, with an approximate sample size of no more than $N=120$ for any dialogue cycle, is statistically insufficient for generalization and, importantly, comprised exclusively of civil society peacebuilders.
Third, while some of the identified areas of common ground may appear self-evident in retrospect, particularly to those engaged in peacebuilding, it is essential to recall the initial context in which this project was undertaken. At that time, a prevailing sentiment held that finding meaningful common ground was unlikely, and the prospect of a constructive joint dialogue was viewed with considerable skepticism.
@@ 4.1. Uninational Phase
Table 2. Coarse-grained shared values and visions emerging from collective statements entailing values produced during the uninational phase. Note that the succinct shared value or vision expressed in each row does not fully entail all ideas contained the associated statements and only captures their abstract common thread. Similarly, presenting statements as associated with the same shared value or vision does not imply equivalency between them.
4.1.1   Red lines
In our work, the term “red lines” refers to actions (including speech acts) that those participating in the process believe are unacceptable, and if undertaken may prevent further dialogue. We asked about two kinds of red lines:
  • Ingroup: the lines that each side themselves would not cross.
  • Outgroup: the lines that each side would not accept the “other side” crossing.
@@ 4.1.1   Red lines
These red lines served a few purposes. Ingroup red lines would provide a trust building signal if they articulated positions that demonstrated respect for the concerns of the other side. The combination of both types of red lines, across all sides, would help clarify the boundary conditions under which a productive joint dialogue could take place. In this context, one notable result of the uninational phase was the mutual alignment between the ingroup and outgroup red lines of Israeli Jews and Palestinians from the West Bank and Gaza. For each of the top three outgroup red lines on each side, the other side had an arguably related ingroup red line that demonstrated some degree of relevant respect and aligned ethics (Table 1). For example, “call for forced displacement” was a top outgroup red line from Palestinians from the West Bank and Gaza, while “talk about deporting Palestinians” was a top ingroup red line from Israeli Jews.
@@
Another notable result came during the transitional deliberation at the end of first dialogue cycle with Israeli Jews. While reviewing one of their top outgroup red lines – “the justification for the attack on October 7th” – strong concern emerged that prohibiting explanation about why October 7th happened would be an unfair constraint on joint dialogue. In the exchange that followed, a mutual understanding emerged: “justification” specifically involved a claim of moral rightness, while an “explanation” was a neutral analysis of causation. So claiming the attack was morally right was the red line, but explaining what caused it to happen was not. This distinction not only applied to many collective statements throughout the process, but helped introduce language like “I won’t justify [X], but an explanation is…” as a rhetorical device for discussing tough issues.
4.1.2   Values and visions
@@
Values and visions were intended to help each side articulate the moral values they held and the future outcomes they sought to achieve. In the spirit of a ”GRIT” (Graduated and Reciprocated Initiatives in Tension reduction) approach (lindskold1978trust), our hypothesis was that commonalities would emerge across all sides that could serve as early points of common ground, prior to the joint dialogue process actually taking place. After all three uninational cycles had concluded, the collective statements entailing values and vision were compared and a set of coarse-grained shared values and visions (Table 2) were identified: peace, equality, human life, independence, safety, and prosperity.
@@ 4.2. Joint phase
Table 3. Collective statements generated during the joint phase along with final vote results, both (a) demands addressed to world leaders, and (b) statements addressed to residents in the region.
@@
Around 120 people participated in the joint phase. Five collective statements comprising joint demands to world leaders were produced (Table 3a). Each statement had at least 90% agreement among participants from each side (i.e., $a_{\textsf{min}}(s,\mathcal{G})>90\%$), and one demand was universally ranked first: “An immediate ceasefire and release of all hostages.” These demands were then composed into a “joint letter” for inclusion in briefs and advocacy efforts:
“We write as a united body of Israeli and Palestinian peacebuilders, compelled by the urgency of our shared pain and galvanized with the hope we still hold for the future. We call for the following: An immediate ceasefire and release of all hostages The creation of an international fund to build the foundations for a comprehensive long-term solution The establishment of a globalized coalition to assist Gaza’s reconstruction The inclusion of civil society and NGOs in all peacebuilding and diplomatic discussions The recognition of both sides’ trauma Acknowledging the trauma and pain experienced by both sides is essential for the healing of our societies and the achievement of sustainable peace. Help us break the cycle of violence so the next generation may flourish. The world watches, and history will judge. Now is the time to act.”
Additionally, five collective statements comprising joint messages to residents in the region were produced (Table 3b). Each of these statements had at least $84\%$ agreement among participants from each side, with most statements being preferred about equally.

### v3

@@ 3.1. Common ground process
3.1.1. Step 1: Collective dialogue
The first step towards a collective dialogue is drafting the planned prompts participants will respond to and deciding what demographics to gather. Prompt drafting is informed by the goals of the dialogue: specific issues to find common ground on, areas to create more mutual understanding around, and so on. Draft prompts are reviewed and iterated until deemed acceptable by a stakeholder group with representation from all sides. The choice of demographics are determined by (a) demographic dimensions likely to manifest relevant divisions, (b) demographics needed to assess representativeness, and (c) constraints imposed by the need for statistical anonymity (Bravo-Hermsdorff et al., 2022).
@@ 3.1.1. Step 1: Collective dialogue
Each collective dialogue generates a a set of natural language statements $S$ responding to each prompt, a sparse sampling of votes on each statement $s\in S$, participant demographics, and data for the non-collective-response questions asked during the dialogue. Vote inference (Konya et al., 2022) is then used to complete the agreement-vote matrix with elements $v(s,p)\in\{1,0\}\equiv\{agree,disagree\}$ that indicate the (stated or inferred) agreement of participant $p$ with statement $s$.
3.1.2. Step 2: Identifying bridging statements
@@
The goal of this step is to identify a set of “bridging” statements $B\subseteq S$ that are likely to represent points of common ground. To do this, we computed two bridging metrics for each statement $s\in S$: (1) max-min agreement $a_{\textsf{min}}(s,\mathcal{G})$ across defined demographic groups $\mathcal{G}$, and (2) the learned parameters $(f_{s},i_{s})$ from the Community Notes algorithm (see Section 2.2.2). As a first set of bridging statements, we selected statements $s\in S$ where $a_{\textsf{min}}(s,\mathcal{G})$ was greater than a given threshold. This gives the set of statements that have the highest lowest agreement across the specified set $\mathcal{G}$ of demographic groups. However, when $G$ contains many demographic groups — defined by age, gender, religiosity, etc. — this approach may exclude a statement that may otherwise bridge a critical axis of division just because it is predicted (perhaps incorrectly) to have low agreement among one demographic sub-group. To mitigate this risk, we created a second set of bridging statements using the Community Notes approach by selecting those where the Euclidean distance $d((f_{s},i_{s}),(0,2))$ was below a given threshold. This approach relies primarily on the note intercept $i_{s}$ as a bridging metric, but also penalizes statements for having non-zero slant, as indicated by the note factor $f_{s}$. The complete set of bridging statements $B$ is taken to be the union of the sets of statements emerging from both approaches.
3.1.3. Step 3: Distillation into collective statements
@@
The set of bridging statements identified in the previous step typically includes multiple restatements of the same idea, compound statements covering multiple ideas, and sub-optimal articulations of ideas. The goal of Step 3 is to create well-articulated versions of each of unique idea that appears in the bridging statements in a way that preserves participants’ choices of specific words and phrases. We use an LLM to do this in two steps. First, the LLM is prompted to extract the unique ideas from the set of bridging statements. Next, the LLM is prompted to generate well-articulated versions of each unique idea in a way that preserves the words and phrases found in the original statements. Included in this prompt are the set of unique ideas, the text of the bridging statements the unique ideas were derived from, and a set of exemplar statements that demonstrate the style in which collective statements should be written. Finally, the collective statements proposed by the LLM are reviewed by local language experts for linguistic consistency with the original bridging statements, and iterated as needed until a final set of collective statements $C$ is deemed fair by a stakeholder group with representation from all sides. The LLM prompts used for this step are included in Appendix H.
3.1.4. Step 4: Final vote
@@
3.1.5. Joint reviews
Even though the final vote serves as a hedge against biases that could cause collective statements to not represent common ground, unfair bias at any point in the process can still undermine legitimacy. To mitigate this risk, at key junctures in the process where bias might be introduced, joint reviews were conducted by representatives of each side before moving on to the next step (see Figure 1). Specifically, joint reviews were conducted on: the content and wording of all messages, questions, and collective response prompts that would be seen by participants in the collective dialogue, the collective statements that would go on to the final vote, and the results computed from the final vote data.
@@ 3.2.2. Strategic sequencing
During the uninational phase, the group that went first needed to do so without the benefit of any trust-building feedback from earlier cycles. Given the existing power asymmetry, it was decided that Israeli Jews would go first. As the second major group, Palestinians from the West Bank & Gaza went second, followed by Palestinian citizens of Israel. Finally, after these three uninational dialogues, we conducted a joint dialogue with participants from all groups to attempt to identify cross-border common ground (Figure 1).
3.3. Participants and Experts
Dialogue participants were recruited via invitations circulated to ALLMEP member organizations (see, e.g. Appendix F). ALLMEP leadership also personally contacted leaders of different organizations to invite participation. Because participants were anonymous and there were multiple stages within each dialogue cycle with no way of tracking identities across stages, it’s not possible to know the exact number of unique participants. We estimate a lower bounds on the number of unique, meaningfully engaged participants in each of the four dialogue cycles to be 70, 28, 13, and 138, respectively (see Appendix G for a detailed breakdown). Additionally, the process included “human experts” in-the-loop, mentioned throughout this Methods section. This small group of experts was hand selected by process facilitators. They consisted of ALLMEP staff and two academics with relevant peacebuilding expertise.
@@ 3.4. Languages
It was critical for legitimacy and inclusivity that participants could fully engage in their primary languages. The uninational cycles would be mono-lingual (Hebrew or Palestinian Arabic), and the joint cycle would need to be multilingual. The main challenge for the mono-lingual cycles was the LLM pipeline used to generate collective statements. Early versions of the pipeline used in prior work (Konya et al., 2023, 2024) had proven effective when the prompts, examples, bridging statements, and generated outputs were all in English, but it was unclear how to approach Hebrew and Palestinian Arabic. Should prompts or examples still be in English? And who decides? To pragmatically address these questions we set up two micro-experiments. We compiled a set of plausible bridging statements in each language and asked native speakers with prompt engineering experience to translate the English prompts and examples. Different variations of the pipeline were created by swapping out different permutations of the English prompts and examples with translated versions. Collective statements were generated using each variation. Then, blind to how they were generated, native speakers were asked which set of collective statements best reflected the content, culture, and language of the original statements they were based on. In all cases, reviewers preferred the statements generated by a pipeline with prompts and examples in their native language, so these variants were used for the uninational phase.
@@ 3.5. Equal power metrics
For the joint cycle, it was critical that result metrics computed from final vote data (ranking and agreement votes) give each side equal influence even with unequal shares of participants. To that end, we computed three equal-power metrics — metrics that are invariant to the portion of participants from each side — for each collective statement: max-min agreement, Dowdall score, and IRV rank.
3.5.1. Max-min agreement
Max-min agreement, the same metric used for bridging-based ranking in Step 2, was computed again using the final agreement votes. Specifically, for each collective statement $s$, max-min agreement was computed as $\min\{\alpha(s,G)\mid G\in\mathcal{G}\}$ where $\alpha(s,G)$ is the fraction of participants in group $G$ that voted “agree” or “strongly agree” on statement $s$. This metric inherently provides equal power across groups and offers a straightforward interpretation: it indicates the minimum level of agreement across all defined groups for statement $s$.
Table 2. Coarse-grained shared values and visions emerging from collective statements produced during the uninational phase. Note that the succinct shared value or vision expressed in each row does not fully entail all ideas contained the associated statements and only captures their abstract common thread. Similarly, presenting statements as associated with the same shared value or vision does not imply equivalency between them.
Table 3. Collective statements generated during the joint phase along with final vote results.
3.5.2. Dowdall score
Similar to a Borda count (Brandt et al., 2016), a Dowdall score is a cardinal metric computed from rank votes using a positional scoring rule that quantifies the strength of relative preference among a set of alternatives for a voting population (Fraenkel and Grofman, 2014). Let the fraction of participants in population $P$ who rank statement $s$ in position $k$ be denoted $\textsf{Position}(s,k;P)$. A Dowdall score is computed as
@@ 3.5.2. Dowdall score
where $\textsf{Points}(k)=1/k$ corresponds to the “points” assigned to each position. We adapt this to be an equal power metric by taking the average of Dowdall scores computed independently for each side:
@@
3.5.3. IRV rank
An instant runoff voting (IRV) rank is an ordinal metric derived from rank votes by iteratively eliminating the option with the fewest first-place votes among remaining options (Brandt et al., 2016). Let $S_{n}$ be the set of statements remaining in round $n$, and $\textsf{First}(s,n;P)$ give the fraction of participants (in some population $P$) who rank statement $s$ first among $S_{n}$. The statements remaining in round $n+1$ are $S_{n+1}=S_{n}\setminus\{\arg\min_{s\in S_{n}}\textsf{First}(s,n;P)\}$. [^5] This formulation assumes no ties for last place, which was true with our data. Denoting $n_{s}$ to be the round where statement $s$ was eliminated, and $|C|$ the total number of statements being ranked, the IRV rank is $\textsf{IRV}(s)=|C|-n_{s}+1$. To make this an equal power metric, the “first place fraction” $\textsf{First}(s,n)$ used to eliminate statements in each round is computed as an average of the first place fractions within each group. That is,
@@ 4. Results
When interpreting the results presented here, it’s important to consider a few contextual factors. First, given the sensitive nature of the issues addressed in these dialogues and their direct relevance to ongoing peacebuilding efforts, the results shared are intentionally selective. Specifically, we present: (1) findings from the uninational phase that demonstrably facilitated the subsequent joint phase, and (2) collective statements of common ground from the joint phase that were explicitly intended for external communication. Furthermore, the quantitative results we share are limited to metrics that give equal power representation across primary participant groups. Second, these findings should not be understood as the views of the broader Israeli and Palestinian populations. The participant pool, with an approximate sample size of no more than $N=138$ for any dialogue cycle, is not only statistically insufficient for generalization, but was comprised exclusively of civil society peacebuilders. Third, while some of the identified areas of common ground may appear self-evident in retrospect, at that time this project us undertaken, a prevailing sentiment held that finding common ground was unlikely, and the prospect of a constructive joint dialogue was viewed with considerable skepticism.
@@ 4.1. Uninational Phase
4.1.1. Red lines
@@
In our work, the term “red lines” refers to actions (including speech acts) that those participating in the process believe are unacceptable, and if undertaken may prevent further dialogue. We asked about two kinds of red lines: ingroup (the lines that each side themselves would not cross) and outgroup (the lines that each side would not accept the “other side” crossing). These red lines served a few purposes. Ingroup red lines would provide a trust building signal if they articulated positions that demonstrated respect for the concerns of the other side. The combination of both types of red line across groups would help clarify the boundary conditions under which a productive joint dialogue could take place. In this context, one notable result of this phase was the mutual alignment between the ingroup and outgroup red lines of Israeli Jews and Palestinians from the West Bank & Gaza. For each of the top three outgroup red lines on each side, the other side had an arguably related ingroup red line that demonstrated some degree of relevant respect and aligned ethics (Table 1). For example, “call for forced displacement” was a top outgroup red line from Palestinians from the West Bank & Gaza, while “talk about deporting Palestinians” was a top ingroup red line from Israeli Jews.
@@
Another notable result came during the transitional deliberation at the end of first dialogue cycle with Israeli Jews. While reviewing a top outgroup red line – “the justification for the attack on October 7th” – strong concern emerged that prohibiting explanation about why October 7th happened would be an unfair constraint on joint dialogue. In the exchange that followed, a mutual understanding emerged: “justification” specifically involved a claim of moral rightness, while an “explanation” was a neutral analysis of causation. So claiming the attack was morally right was the red line, but explaining what caused it to happen was not. This distinction not only applied to many collective statements throughout the process, but helped introduce language like “I won’t justify [X], but an explanation is…” as a rhetorical device for discussing tough issues.
4.1.2. Values and visions
@@
Values and visions were intended to help each side articulate the moral values they held and the future outcomes they sought to achieve. In the spirit of a “GRIT” (Graduated and Reciprocated Initiatives in Tension reduction) approach (Lindskold, 1978), our hypothesis was that commonalities would emerge across all sides that could serve as early points of common ground, prior to the joint dialogue process actually taking place. After all three uninational cycles had concluded, the collective statements entailing values and vision were compared and a set of coarse-grained shared values and visions (Table 2) were identified: peace, equality, human life, independence, safety, and prosperity.
@@ 4.2. Joint phase
Around 138 people participated in the joint phase. Five collective statements comprising joint demands to world leaders were produced (Table 3a). Each statement had at least 90% agreement among participants from each side (i.e., $a_{\textsf{min}}(s,\mathcal{G})>90\%$), and one demand was universally ranked first: “An immediate ceasefire and release of all hostages.” These demands were then composed into a “joint letter” for inclusion in briefs and advocacy efforts (Appendix A). Additionally, five collective statements comprising joint messages to residents in the region were produced (Table 3b). Each of these statements had at least $84\%$ agreement among participants from each side, with most statements being preferred about equally.